OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (1 point) 5x – y = –1 7x – 3y = 13

OpenStudy (radar):

I can explain how I would use substitution, but keep in mind there are other ways. First I would use the first equation to isolate y.

OpenStudy (anonymous):

u dont need to use substitution to solve this problem!

OpenStudy (anonymous):

I want to try and use substitution.

OpenStudy (anonymous):

I know elimination isbetter

OpenStudy (radar):

this is accomplished as follows. Add y to both sides of the equal sign of the first equation getting: y-1=5x. I would then add 1 to both sides of this equation getting: y=5x+1. this has now isolated y and its value is substituted for the y in the second equation. Follow so far?

OpenStudy (anonymous):

yes, but i just learned this so im a little rough, can you continue

OpenStudy (anonymous):

you have to isolate the y right?

OpenStudy (radar):

Yes now substituting for the y in the second equation. Remember that we have already determined that y=5x+1. When substituted for y in the second equation we would now have: 7x-3y=13 or 7x-3(5x+1)=13 Do you see where y was placed in the term -3y ?

OpenStudy (radar):

Please review my second post to determine how we isolated y.

OpenStudy (anonymous):

Uhhh i think so. but i dont kow how to finish it all off. Do i just cancel everything?

OpenStudy (radar):

By substituting the isolated value for y in the second equation, we now have one equation in one unknown. That is 7x-15x-3=13. All you have to do is to read and understand these posts.

OpenStudy (anonymous):

Ohh oh i c now.

OpenStudy (radar):

Do you know how to solve 7x-15x-3=13 ?? Using the common principles of algebra.

OpenStudy (anonymous):

yuppp. that helps

OpenStudy (radar):

Just remember the Golden Rule, " what ones does to one side of the equal sign, one must do the same thing to the other side"

OpenStudy (radar):

I think you have it but I will continue. Combine the x terms getting: -8x-3=13 Now add 3 to both sides getting -8x=16 Divide both sides by -8 getting x=-2 You now have x. Substitute this in one of the original equations (your choice of which one) and then solve for y.

OpenStudy (anonymous):

Ok, thanks bro!

OpenStudy (anonymous):

yes! thats what i got for x and now y is simple.

OpenStudy (anonymous):

did you get y=4?

OpenStudy (radar):

You're welcome and you are now on your way. good luck with these system of linear equations.

OpenStudy (anonymous):

yeah?

OpenStudy (radar):

Let me check 5x-y=-1 5(-2)-y=-1 -10-y=-1 Doesn't look like I am going to get the same thing as you did.

OpenStudy (radar):

Double check you work, pay real close attention to your signs!

OpenStudy (anonymous):

hmmmm. ill recheck my math. thanks for the help!

OpenStudy (radar):

Which equation did you choose to substitute the -2 for x?

OpenStudy (anonymous):

7x-3y=13

OpenStudy (radar):

So substituting -2 for x that would look like: 7(-2) - 3y = 13 -14 -3y = 13 adding +14 to both sides gets us -3y = 27 dividing by both sides by -3 get us y = 9 Where did you stray off the "straight and narrow"?

OpenStudy (radar):

That should be a -9 please excuse the typo?

OpenStudy (radar):

See how easy it is to get signs confused!

OpenStudy (radar):

Earth to @NashNugs are you breathing?

OpenStudy (radar):

I have tried to comply with your problem, the sentences were supposed to be complete.

OpenStudy (anonymous):

Ooops sorry! I was just trying to figure out the proble. I got -9 i see where i goofed up!

OpenStudy (radar):

take care. Where did you goof up? Just curious.

OpenStudy (anonymous):

Thanks! Oh i substituted wrong.

OpenStudy (anonymous):

i gotta go! i appreciate it sir!